OpenStudy (anonymous):

Which of the following is the equation of a parabola with focus (0, 2) and directrix y = -2? a. x=1/8y^2 b.y=1/8x^2 c.y=1/4x^2

OpenStudy (thatonegirl_):

\[x^2=4(p)y\]

OpenStudy (thatonegirl_):

this is your formula because the directrix is always the opposite of the squared value (or the one that starts the equation). Since it is y=-2, its gonna be x^2.

OpenStudy (thatonegirl_):

p is going to be 2, because that is the other value in the focus. \[x^2=8y\]

OpenStudy (thatonegirl_):

\[1x^2=8y\]

OpenStudy (thatonegirl_):

\[\frac{ 1 }{ 8 }x^2=y\]

OpenStudy (jdoe0001):